hi, my name is lee. thank you for helping me
I would like to know how to do this word problem.
Three times and odd interger is eleven less then four times the next greatest even integer. What is the odd interger?
could you please show me the formula?
thank you
lee

Answer by josmiceli(19441)

K = the odd integer
K + 1 = the even integer
3K = 4(K + 1) - 11
3K = 4K + 4 - 11
K = 11 - 4
K = 7
K + 1 = 8
3(7) = 4(8) - 11 check results
21 = 32 - 11